OpenStudy (nightmarenight):

Will give medal for right answer!! y varies inversely with x and y = 8 when x = 2. What is the inverse variation equation for the relationship?

zepdrix (zepdrix):

`y varies inversely with x` \[\Large\bf\sf y=\frac{1}{x}\]This relationship shows that when we have a y value, x is the multiplicative inverse of y. (Reciprocal). But we also need a constant of variation.\[\Large\bf\sf y=\frac{k}{x}\]

zepdrix (zepdrix):

So they gave us some information to work with. y=8 when x=2. We'll plug in this relationship to solve for our constant k.

zepdrix (zepdrix):

Understand how to plug them in ? :O

OpenStudy (nightmarenight):

so 8=k/2 ??? @zepdrix

zepdrix (zepdrix):

Mmmm k good! So k = ?

OpenStudy (nightmarenight):

16? @zepdrix

zepdrix (zepdrix):

Mmm good! Plug that back in your expression with the x and y and you have your answer.

OpenStudy (nightmarenight):

Wait 16 isn't the answer? im confused now /: @zepdrix

zepdrix (zepdrix):

Here is the relationship we have, and here is what you found out about k. \[\Large\bf\sf y=\frac{\color{royalblue}{k}}{x}, \qquad\qquad\qquad \color{royalblue}{k=16}\]

zepdrix (zepdrix):

Plug it in!! :O Plug in the blue!

OpenStudy (nightmarenight):

8= 16/2 ?? @zepdrix

zepdrix (zepdrix):

Ok you have the right idea, but you no longer want the x=2, y=8 plugged in. Only the k. That gives us our relationship for x and y.\[\Large\bf\sf y=\frac{\color{royalblue}{16}}{x}\]